OpenStudy (anonymous):

What is the area and perimeter of a rectangle who's side length is 5cm and width is 3cm?

OpenStudy (anonymous):

im not sure, but i guess either 15, or 16. either or, pretty sure.

OpenStudy (anonymous):

cuz 5*3=15 or 5*2=10+3*2=6=16. maybe.

OpenStudy (anonymous):